OpenStudy (anonymous):

i have to write and equation in slope inercept form for the line that passes throughthe given point and is parallel to the graph equation (5,4), y=2/5x - 2 (3,1), 2x + y=5 (-3,4), 3y=2x-3 (-,-2), 3x - y=5 (-8,2), 5x - 4y=1 (-1,-4), 9x + 3y=8 (-5,6), 4x + 3y=1

hero (hero):

That's a lot of equations

OpenStudy (anonymous):

there were 24 on the hole page and those are the one section on the page i couldnt figure out the rest was grapging n stuff i got that :/

hero (hero):

Oh wow. What do you already know about this?

OpenStudy (anonymous):

i need to find y n i for get how to do ppl have showed me different ways but they look different and they all did it different so im confused how to do them

hero (hero):

The first one is already solved for y

hero (hero):

Do you know how to proceed from there?

OpenStudy (anonymous):

i was told the first one was wen u find the y inercept that its y=2/5x+2 i have to find where they pass through like that i have the fisrt one someone gave me that answer

hero (hero):

I see. But you should have just posted the original equation that was given. How would I know that the first one is correct?

OpenStudy (anonymous):

those are the original equastions

hero (hero):

You're not understanding me. The first answer that someone gave you....there's no guarantee that it is correct.

hero (hero):

Are you saying that for the first one, what you have posted is the original equation?

OpenStudy (anonymous):

ohh okay so what do i do

OpenStudy (anonymous):

yes there all the originall questions

hero (hero):

Basically, the instructions says that the line must be parallel to the given line and it must pass through the given point. So 1. Find the slope of the line parallel to the given line, 2. Use y = mx + b to find the y-intercept. 3. Rewrite the equation in the form y = mx + b but include the values for m and b.

OpenStudy (anonymous):

no i forget how

OpenStudy (anonymous):

So you were given the correct answer to equation #1. Right? So if I understand you, then what you are trying to find is a new b value using the point given. Changing the b value will shift the line up or down.

hero (hero):

We can go on vyew if you like...I can probably explain it better there.

OpenStudy (anonymous):

ya can u help me with them plzz it feels wierd im stressin over this and its not even mine i figured id get it it since i already did it lol

hero (hero):

Whosever it is, you should give it back to them and make them do it.

hero (hero):

I can explain it better on vyew.

OpenStudy (anonymous):

my lil sisters shes failing math so i told her id so it and figure it out so i can help her

OpenStudy (anonymous):

Here's a graph for the second equation.

OpenStudy (anonymous):

i dont have to graph it i just need the problem thingy

hero (hero):

Are you ready to go on vyew or not?

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

how do i get there

hero (hero):

When you go there, make sure you identify yourself using your OS name only

OpenStudy (anonymous):

To find a parallel line to the equation given and the point given all you need to do is find the new b value. So for equation 2: rearrange 2x +y= 5 to -2x+5=y Now set the number 5 to b and plug in the given point (3,1) for the x and y values So now you have (-2)(3) + b = 1 Solve for b. So now you have b =1+6 which is 7. Your new equation is now -2x+7=y

OpenStudy (anonymous):

thank u @marshallinwashington can u hel pme with more plz :)

OpenStudy (anonymous):

I'll do one more and then you try to do rest on your own. If you get stuck we can help you.

hero (hero):

@marshallinwashington , you were supposed to find the slope of the parallel line first, not just put it in the form y = mx+b

OpenStudy (anonymous):

For equation 3y=2x-3, rearrange the equation by dividing both sides by 3. So now it looks like this: y=2/3x - 3/3 which becomes y=2/3x - 1 Replace 1 with b and plug in the point given (-3,4) So now you have 4=(2/3)(-3)-b That becomes 4= -2-b Add 2 to both sides. Which becomes 6=-b Multiply both sides by -1 So b = 6 So your new equation becomes y=2/3x +6

hero (hero):

@marshallinwashington , did you read the part of the instructions that said to find the line PARALLEL to the given line?

OpenStudy (anonymous):

Yes. A new parallel line equation will have the same slope as the original equation. The only thing that needs to be done is to find the new b value for the given point. That's the way I'm reading the question.

hero (hero):

Yes, you are right. My bad...I was thinking perpendicular lines.

hero (hero):

Now I have to erase everything I wrote

OpenStudy (anonymous):

Ok, now that we both agree, We can give mak_12 a hand if needed.

hero (hero):

Well, I think she's gone already. I was helping her on vyew before and she said something about softball practice.